OpenStudy (anonymous):

Solve 5/(2x+1)?? please show your work.

OpenStudy (anonymous):

What are we solving for here?

OpenStudy (anonymous):

x i believe.

OpenStudy (anonymous):

What does it equal?

OpenStudy (anonymous):

my text says something about the domain of a variable. if that helps

OpenStudy (anonymous):

i got 10. solve 2x-1 = 0 you get 1/2 or 0.5 5 divided by 0.5 would be 10. not sure if correct, but it was a shot.

OpenStudy (anonymous):

could be -10 as well 5/2x - 1 multiply both numerator and denominator by 2 give: 10/-1 which equals -10....just throwing ideas, hopefully someone can check me on my math.

OpenStudy (anonymous):

nvm that would leave the denominator as x-1, i screwed that up....

OpenStudy (anonymous):

by book gave the answer 1/2

OpenStudy (anonymous):

Okay. The domain is the range of values for a variable where a function is defined. \[\text {Let}~~ f(x) = {5 \over (2x+1)}\]A rational function (rational means fraction) is undefined when the denominator is equal to zero. Therefore, \[2x+1 \ne 0\]The denominator is equal to zero only when x = -1/2. Therefore, the domain, in set notation, is \[(-\infty,-{1 \over 2}) \cup (-{1 \over 2}, + \infty)\]

OpenStudy (anonymous):

its 2.5 2x+1 dividing by 2 gives x=1/2 or 0.5 thus, 2(0.5) =1 1+1 =2 5/2 =2.5 im sure of this.

OpenStudy (anonymous):

its another way of saying 1/2 so hollywood_chrissy was right.

OpenStudy (anonymous):

you mean eastmore was right

OpenStudy (anonymous):

eashmore I believe its more of an algebraic problem than it is funstion and set notation, otherwise it would have been stated as such. @ Hollywood: sorry meant your book was right.